I need help setting this problem up please... The question reads: "Two trains, an express and a commuter , are 150 miles apart. Both start at the same time and travel toward each other. They meet 3 hours later. The speed of the express is 10 miles faster than the commuter. Find the speed of each train." Thanks for you help!!
Answer by ptaylor(2198)


ok
Distance(d) equals Rate(r) times Time(t)or d=rt; r=d/t and t=d/r
Let r=speed of the commuter
Then r+10=speed of the express
distance commuter travels=rt=r*3 or 3r
distance express travels=rt=(r+10)*3 or 3r+30
Now we are told that the above two distances add up to 150 mi, so:
3r+3r+30=150 Can you handle it from here???

Hope this helps---ptaylor